What is the product of the complex # (3+i) and (3-i)
Found 2 solutions by ichudov, Alan3354:
Answer by ichudov(507)

Solved by pluggable solver: ARITHMETIC on Complex Numbers
6 +0i.


Solution:Addition of two complex numbers a + bi
and c+di is done according to formula:



(a + bi) + (c+di) = (a+c) + (b+d)i.


In our case, .

Answer by Alan3354(69443)

What is the product of the complex # (3+i) and (3-i)
----------------------
= 3^2 + i^2 + 3i - 3i
= 9-1
= 8